find dy/dx y= (e)^x + (10)^x + (x)^x

find dy/dx
y= (e)^x + (10)^x + (x)^x

Grade:10

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\\frac{d}{dx}\left(e^x+10^x+x^x\right)=e^x+10^x\ln \left(10\right)+x^x\left(\ln \left(x\right)+1\right) \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty